OpenStudy (anonymous):

carson drives to school the same way each day and there are two independent traffic lights on his trip to school. He knows that there is a 30% chance that he will have to stop at the first light and an 80% chance that he will have to stop at the second light. What is the probability that he will NOT have to stop at either light? 14% 24% 50% 80%

OpenStudy (anonymous):

Okay Yayo. So 30% and 80% chance that he will have to stop at one of the lights. So what we want to find is the difference of that percentage from 100% so we know the probability of him NOT having to stop.

OpenStudy (anonymous):

100 - 30 = 70% and 100-80 = 20%

OpenStudy (anonymous):

So let's multiply them together to find the probability. 7/10 * 1*5. What do you get?

OpenStudy (anonymous):

14%?

OpenStudy (anonymous):

A two-way frequency table shows grades for students in college and students in high school.

OpenStudy (anonymous):

Based on this data, are "being in high school" and "GPA above 3.0" independent events? Yes, P(high school | GPA above 3.0) = P(high school) Yes, P(high school | GPA above 3.0) = P(GPA above 3.0) No, P(high school | GPA above 3.0) ≠ P(high school) No, P(high school | GPA above 3.0) ≠ P(GPA above 3.0)

OpenStudy (anonymous):

@Lethal

OpenStudy (anonymous):

Sorry. Yes I'm back. It's 14% correct.

OpenStudy (anonymous):

For the second one. I'm not sure. Try posting it in a new question :)